If a penny is 19.5 mm in diameter, how many pennies are needed to stretch from the earth to the sun? It takes 8.00 minutes and 30.0 seconds for light to travel from the sun to the earth at a speed of 186,282 miles per second.
1
Expert's answer
2012-09-14T10:06:59-0400
The distance from the earth to the sun is l=c*t=299,792,458 metres per second *
( 8*60 + 30 ) seconds = 152894153580 metres .
The diameter of a penny is
0.0195 metres.
Then we need 152894153580/0.0195=7840725824616 pennies